### v1

# On the extension of a class of Hermite bivariate interpolation problems

Abstract
We characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space. This result extends an earlier theorem (2000) by one of the authors concerning the case . The latter theorem, in turn, can be regarded as a natural generalization of a classical theorem of Severi (1921).
Key words: Fundamental polynomial, algebraic curve, -correct set, -solvable set.
Mathematics Subject Classification (2010):
primary: 41A05, 41A63; secondary 14H50.
1 Introduction
In this paper, we characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1). This extends a previous result (Theorem 1.20) that addresses the case of . The latter theorem can, in turn, be regarded as a natural generalization of a classical theorem of Severi (Theorem 1.21).
One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M. Nagata [6]. In approximation theory the investigation was initiated by G. G. Lorentz and R. A. Lorentz [4].
It should be noted also that a similar study in the case of Lagrange bivariate interpolation is more advanced (see [1]).
1.1 Lagrange bivariate interpolation
Denote the space of bivariate polynomials of total degree with real or complex coefficients:
Let us fix a set of distinct nodes in the plane:
The problem of finding a polynomial satisfying the conditions
for a given data is called Lagrange interpolation problem.
Definition 1.1.
A set of nodes is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying the conditions (1.2).
Denote by the restriction of on the set
A polynomial is called an -fundamental polynomial for if
We denote this -fundamental polynomial by
Let us mention that sometimes we call -fundamental a polynomial satisfying the conditions
since it equals a non-zero constant times
Definition 1.2.
A set of nodes is called -independent if each node has an -fundamental polynomial. Otherwise, it is called -dependent.
If a set is -independent then the following Lagrange formula gives a solution to the problem (1.2):
Thus we obtain:
Proposition 1.3.
A node set is -independent if and only if it is -solvable.
Fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, is
Definition 1.4.
A set of nodes is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.2).
A necessary condition of -correctness is:
For the sets of nodes satisfying the above condition we have
Proposition 1.5.
Assume that Then the following conditions are equivalent:
The set is -correct.
The set is -solvable.
A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation . In particular, by we denote a linear polynomial from and the line defined by the equation Similarly by we denote a quadratic polynomial from and the conic defined by the equation
1.2 Hermite bivariate interpolation
Hermite interpolation problem is determined by a pair where is a set of nodes given by (1.1) and is a scheme, i.e., a sequence of multiplicities of the nodes associated with the degree
As a rule we assume that the multiplicities in the scheme are non-increasing:
Denote
In the problem the interpolation parameters at a given node are the value of a polynomial and its partial derivatives up to the order
The problem of finding a polynomial satisfying the above conditions, for a data
is called Hermite interpolation problem.
We express briefly the conditions (1.4) as follows
Nodes whose multiplicity is equal to one are called simple. At these nodes only the value of a polynomial is interpolated in the conditions (1.4).
It is sometimes convenient to allow nodes of multiplicity in the problem In fact, such a node can be ignored because there is no interpolation condition.
Definition 1.6.
A problem is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying (1.4).
A polynomial is called an -fundamental polynomial for a parameter if it is a solution to the problem (1.4) where
The above fundamental polynomial we denote by
Definition 1.7.
A problem is called -independent if for each parameter in (1.4) there is an -fundamental polynomial.
The Lagrange formula for the above problem giving a solution to the problem (1.4), takes the form:
From here we get:
Proposition 1.8.
A problem is -independent if and only if it is -solvable.
Denote
We have that the number of interpolation conditions at a node in (1.4) equals to Hence the total number of the conditions equals to
It is easily seen that fundamental polynomials are linearly independent. Also we have that Therefore a necessary condition of -independence, or -solvability, of the problem is
A scheme satisfying the above condition we call -scheme.
A scheme satisfying the condition
we call -scheme.
Definition 1.9.
A problem is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.4).
A necessary condition of -correctness is:
A scheme satisfying the above condition we call an exact scheme.
Below we have the analog of Proposition 1.5 for the Hermite interpolation.
Proposition 1.10.
Assume that a scheme is exact. Then the following conditions are equivalent:
The problem is -correct.
The problem is -solvable.
Proposition 1.11.
Assume that in the problem the scheme is -scheme. Then there exists
Indeed, note that is the number of interpolation conditions in (1.4) and each condition can be considered as a linear equation with respect to unknown coefficients of Thus we get a linear homogeneous system, where the number of equations is less than the number of unknowns. As we know such systems have non-trivial solutions.
In the simplest case the above proposition implies
Lemma 1.12.
For any five points, there exists a conic that passes through all of them.
For a -scheme we define the exact scheme
Accordingly, for a set we define
where is a set of simple points.
Proposition 1.13 ([5], lemma 2.1).
The problem is -correct, for some if and only if the problem is -solvable.
For the sake of completeness we present:
Proof.
The direction “only if” is evident. For the direction “if”, in view of Proposition 1.10, it is enough to prove that there exists a (simple) node such that the problem where and is -solvable. For this, in view of -solvability of we need just an -fundamental polynomial of the simple node
It is easily seen that the polynomial is the desired -fundamental polynomial of ∎
Let and
Remark 1.14 ([3], rem.1.14).
If the problem is -solvable, then so is the problem with sufficiently close to
Definition 1.15.
Let be a problem, where is given by (1.1), and be a curve of degree Then we define the scheme as follows
where if and otherwise.
Denote by the number of nodes from on the curve counted with the multiplicities from
Note that in what follows, this notation, as well as the previous definition, mainly will be used in cases where the curve is a line or a conic.
We will need the following well-known proposition (see e.g. [2], lemma 4):
Proposition 1.16.
Let be a scheme with and be a set of nodes with where is a line. Then
Set
The following proposition we will use for lines and (irreducible) conics.
Proposition 1.17.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then for any curve of degree we have that
Proof.
Let the problem be -solvable for a set of nodes Assume, by way of contradiction, that for a curve of degree we have that
Then, Proposition 1.13 implies that
for some where is a set of simple points with and
Moreover, by making use of Remark 1.14 we may assume that the set contains no point from the curve
Consider the following scheme
The above condition on the curve in view of (1.7), means that
Then let us check that is -scheme. Indeed, by using (1.7) and (1.9), we get
Therefore, in view of Proposition 1.11, there exists a polynomial such that
Now one readily obtains for the polynomial that
This, in view of Proposition 1.10, contradicts the -correctness of the problem ∎
By applying Proposition 1.17 for the cases we obtain the following two necessary conditions for -solvability:
Corollary 1.18.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then:
For any line we have that
For any conic we have that
In the sequel the condition (i) will be referred to as the -condition.
Similarly, the condition (ii) will be referred to as the -condition.
Note, that by taking into account the above condition (i) for the lines passing through any two nodes, we obtain
Corollary 1.19.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is solvable for a set of nodes then
Now we are in a position to present the result (slightly reformulated), which we will extend.
Theorem 1.20 ([2], th. 9).
Let be a scheme, satisfying the condition
Then the problem is -solvable if and only if there is no line passing through more than nodes of
This theorem can be regarded as a natural generalization of the following classical theorem of Severi (1921).
Theorem 1.21 (Severi [7], th. 9).
Let be a scheme. The problem is solvable for any node set if and only if
From Theorem 1.20 we get
Corollary 1.22.
For any scheme with non-increasing sequence of multiplicities, we have that
Indeed, consider the problem where in no three nodes are collinear. Then according to Theorem 1.20 the problem is -solvable. Hence is -scheme.
Proposition 1.23.
Let be a scheme with non-increasing sequence of multiplicities such that
Suppose is a set of nodes with where is an irreducible conic. Then we have that
Proof.
Consider the following scheme:
We have that
Note that
The above inequality, in view of irreducibility of implies that the -condition is satisfied. Thus, in view of Theorem (1.20), we obtain that the problem is -solvable. Similarly we obtain that the problem is -solvable.
Now consider the following linear spaces:
Note that In view of … we have that
Thus we obtain that ∎
Below we consider the case of above proposition. Note that the cases are not possible.
Remark 1.24.
Let be a scheme with non-increasing sequence of multiplicities such that
Then we readily get that implies that While implies that
Both these schemes are exact. According to the above proof the problems are also solvable, provided with any irreducible conic. The latter condition in this case means that no three nodes are collinear. Thus the problems are correct:
This holds also with the scheme replaced by Therefore Proposition 1.23 is valid also in the case since the statement reduces to
Remark 1.25.
Note that according to Corollary 1.22 for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme. Indeed, for a counterexample, we just add a fifth multiplicity to the schemes and of the above remark.
This is why in the next theorem, unlike Theorem 1.20, we have the condition: is -scheme.
2 The result
In this section we will prove the following
Theorem 2.1.
Let be a -scheme satisfying the condition
Then the problem is -solvable if and only if
there is no line passing through more than nodes of ,
there is no conic passing through more than nodes of
Proof.
The direction “only if” follows from Corollary 1.18. Let us prove the direction “if”.
Note that the above condition (ii) means that there is no conic passing through all the nodes of
In view of Theorem 1.20 we can assume, in what follows, that
From now on suppose also that is a scheme with non-increasing sequence of multiplicities:
In view of Corollary 1.19 we have that
We can assume also that
Indeed, the only - scheme, in the case is the Taylor scheme which does not satisfy (2.2). While if then the conditions (2.2) and (2.4) imply that which is not - scheme.
We will apply induction on the degree of the scheme
Step 1. It is easily seen that if then there is no -scheme satisfying the condition (2.2).
In the case there is only one - scheme satisfying the conditions (2.2) and (2.5). According to the condition (ii) of Theorem, the six simple nodes of the set do not belong to a conic. Since is exact scheme, by Proposition 1.5 (iii), we get that is -correct.
Now, assuming the statement holds for all degrees we will prove it for where
Step 2. Assume that there is a line which passes through nodes of i.e., the following condition is satisfied:
Consider the scheme (see (1.6))
From (1.7), in view of the condition (2.6), we have that
Now let us check that is - scheme. Indeed, by using (2.7), we get
The condition (2.1) for readily follows from the fact that passes through at least two nodes of
Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing nodes of Assume conversely that there is such a line Then evidently and, in view of the relation
we obtain that contains nodes from Since contains nodes both from and we conclude that the line does not intersect at a node of Therefore, the (reducible) conic passes through nodes of which contradicts the condition (ii) of Theorem.
Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing nodes of In view of (2.9) this means that passes through all the nodes with
Note that if passes through at least three nodes of then in (2.9) we would obtain and hence the -condition is satisfied. It remains to consider the case where passes through exactly two nodes, say, In this case, in view of (2.6), we have that According to (2.5) we get that Hence we have that Thus passes through all the nodes of which is contradiction.
Therefore, the problem is -solvable by virtue of the induction hypothesis.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
and the nodes of correspond to the last ones in
Here, by making use of Remark 1.14, we can also assume that the set contains no node belonging to the line
Now consider the problem where
Let us check first that is exact scheme. Indeed, in view of (2.7), we have that
Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.10, (iii), that where and prove that
By the choice of the set and the condition (2.6) we have that
Now, Proposition 1.16 implies that where
Now, in view of (2.10) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies evidently that is -solvable. ∎
From now on, in view of Step 2, we can assume that the -condition is satisfied for the scheme which implies that
Step 3. Assume that there is a line passing through at least nodes of Consider the following scheme (see (1.6))
We have that
On the other hand, in view of Step 2, the -condition holds. Thus according to Theorem 1.20 we obtain that the problem is -solvable.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
Again, by making use of Rremark 1.14, we can assume also that the set contains no node belonging to the line
Then denote by a set of simple nodes in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
In view of the choice of the sets and we have that
Therefore, Proposition 1.16 implies that where
Now, in view of (2.13) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎
From now on, in view of Step 3, we may assume that there are no three collinear nodes in the set
Step 4. First note that the condition (ii) of Theorem, in view of Lemma 1.12, implies that
Consider a conic passing through the nodes Since no three nodes are collinear the conic is irreducible.
Note that may not satisfy the condition (2.3). Thus, to apply Theorem 1.20, let us verify that
Indeed, if then and, in view of (2.12), this is evident.
Now, assume conversely, that the inequality holds for some
Then, in view of (2.3), we have that and therefore
On the other hand we have that if Hence we have that which contradicts the condition (2.1).
Thus, according to Theorem 1.20, the problem is -solvable.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
By making use of Remark 1.14 we can assume that the set contains no node belonging to the conic
Next, denote by a set of simple nodes in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
By taking into account the choice of the sets and we have that Therefore, Proposition 1.23 implies that where
Now, in view of (2.15) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎

### v2

@@
-On the extension of a class of Hermite bivariate interpolation problems
+On the extension of a class of Hermite multivariate interpolation problems
 Abstract
-We characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space.
+We characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space.
 This result extends an earlier theorem (2000) by one of the authors concerning the case .
-The latter theorem, in turn, can be regarded as a natural generalization of a classical theorem of Severi (1921).
-Key words: Fundamental polynomial, algebraic curve, -correct set, -solvable set.
+The latter theorem, in turn, can be regarded as a natural extension of a classical Theorem of Severi (1921).
+Key words: Hermite interpolation, fundamental polynomial, algebraic curve, -correct set, -solvable set.
 Mathematics Subject Classification (2010):
@@
 1 Introduction
-In this paper, we characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1).
-This extends a previous result (Theorem 1.20) that addresses the case of .
-The latter theorem can, in turn, be regarded as a natural generalization of a classical theorem of Severi (Theorem 1.21).
+In this paper, we characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1).
+This result extends a previous one (Theorem 1.18) that addresses the case of .
+The latter theorem can, in turn, be regarded as a natural extension of a classical Theorem of Severi (Theorem 1.19).
 One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M.
@@
 A.
-Lorentz [4].
+Lorentz [4], [5].
 It should be noted also that a similar study in the case of Lagrange bivariate interpolation is more advanced (see [1]).
-1.1 Lagrange bivariate interpolation
-Denote the space of bivariate polynomials of total degree with real or complex coefficients:
-Let us fix a set of distinct nodes in the plane:
-The problem of finding a polynomial satisfying the conditions
-for a given data is called Lagrange interpolation problem.
-Definition 1.1.
-A set of nodes is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying the conditions (1.2).
-Denote by the restriction of on the set
-A polynomial is called an -fundamental polynomial for if
-We denote this -fundamental polynomial by
-Let us mention that sometimes we call -fundamental a polynomial satisfying the conditions
-since it equals a non-zero constant times
-Definition 1.2.
-A set of nodes is called -independent if each node has an -fundamental polynomial. Otherwise, it is called -dependent.
-If a set is -independent then the following Lagrange formula gives a solution to the problem (1.2):
-Thus we obtain:
-Proposition 1.3.
-A node set is -independent if and only if it is -solvable.
-Fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, is
-Definition 1.4.
-A set of nodes is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.2).
-A necessary condition of -correctness is:
-For the sets of nodes satisfying the above condition we have
-Proposition 1.5.
-Assume that Then the following conditions are equivalent:
-The set is -correct.
-The set is -solvable.
-A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation .
-In particular, by we denote a linear polynomial from and the line defined by the equation Similarly by we denote a quadratic polynomial from and the conic defined by the equation
-1.2 Hermite bivariate interpolation
-Hermite interpolation problem is determined by a pair where is a set of nodes given by (1.1) and is a scheme, i.e., a sequence of multiplicities of the nodes associated with the degree
+1.1 Hermite multivariate interpolation
+Hermite interpolation by multivariate algebraic polynomials is considered. The interpolation parameters are the values of a function and its partial derivatives up to some orders at given points.
+Denote for and the multi-index
+Set for convenience
+The space of polynomials in variables of total degree is denoted by
+The coordinates of points and coefficients of polynomials can be taken either real or complex, thus leading to a real or complex interpolation problem, respectively.
+Consider a set of distinct interpolation points
+Consider also a scheme, i.e., a sequence of multiplicities of the points associated with the degree of interpolating polynomials
 As a rule we assume that the multiplicities in the scheme are non-increasing:
-Denote
-In the problem the interpolation parameters at a given node are the value of a polynomial and its partial derivatives up to the order
+In Hermite interpolation problem where is the dimension, the interpolation parameters at a given point are the value of a polynomial and its partial derivatives up to the order
 The problem of finding a polynomial satisfying the above conditions, for a data
-is called Hermite interpolation problem.
-We express briefly the conditions (1.4) as follows
-Nodes whose multiplicity is equal to one are called simple. At these nodes only the value of a polynomial is interpolated in the conditions (1.4).
-It is sometimes convenient to allow nodes of multiplicity in the problem In fact, such a node can be ignored because there is no interpolation condition.
+is called Hermite interpolation problem
+From now on we express briefly the conditions (1.4) as follows
+Points whose multiplicity is equal to one are called simple. At these points only the value of a polynomial is interpolated in the conditions (1.4).
+It is sometimes convenient, for example in proofs, to allow points of multiplicity in the problem In fact, such a point can be ignored because there is no interpolation condition.
 Definition 1.1.
@@
 The above fundamental polynomial we denote by
+Let us mention that sometimes we call -fundamental a polynomial, which is a solution to the problem (1.4) where since it equals a non-zero constant times
 Definition 1.2.
-A problem is called -independent if for each parameter in (1.4) there is an -fundamental polynomial.
+A problem is called -independent if for each parameter in (1.4) there exists an -fundamental polynomial.
 The Lagrange formula for the above problem giving a solution to the problem (1.4), takes the form:
@@
 A problem is -independent if and only if it is -solvable.
-Denote
-We have that the number of interpolation conditions at a node in (1.4) equals to Hence the total number of the conditions equals to
-It is easily seen that fundamental polynomials are linearly independent. Also we have that Therefore a necessary condition of -independence, or -solvability, of the problem is
+We have that the number of interpolation conditions at a point in (1.4) equals to (see 1.1). Hence the total number of the conditions equals to
+It is easily seen that fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, of the problem is
 A scheme satisfying the above condition we call -scheme.
@@
 A necessary condition of -correctness is:
-A scheme satisfying the above condition we call an exact scheme.
-Below we have the analog of Proposition 1.5 for the Hermite interpolation.
+A scheme satisfying the above condition we call exact scheme.
 Proposition 1.5.
-Assume that a scheme is exact. Then the following conditions are equivalent:
-The problem is -correct.
+Assume that a scheme is exact. Then the -correctness of the problem is equivalent to each of the following conditions:
 The problem is -solvable.
+Proof.
+Indeed, each of the conditions in (1.4) can be considered as a linear equation with respect to unknown coefficients of Thus for exact scheme the conditions (1.4) reduce to a square linear system (see (1.6)).
+Now, the condition (i) means that the linear system is consistent for any right-hand side values, while the condition (ii) means that the homogeneous linear system has a single trivial solution.
+It is well known that in the case of a square system each of these properties implies that the system has a unique solution for any right-hand side values.
+The later property means that the problem is -correct.
+∎
 Proposition 1.6.
 Assume that in the problem the scheme is -scheme. Then there exists
-Indeed, note that is the number of interpolation conditions in (1.4) and each condition can be considered as a linear equation with respect to unknown coefficients of Thus we get a linear homogeneous system, where the number of equations is less than the number of unknowns.
+Proof.
+Indeed, in this case the conditions (1.4) reduce to a linear homogeneous system, where the number of equations is less than the number of unknowns.
 As we know such systems have non-trivial solutions.
-In the simplest case the above proposition implies
-Lemma 1.12.
-For any five points, there exists a conic that passes through all of them.
-For a -scheme we define the exact scheme
+∎
+Thus, for the -scheme we obtain
+Corollary 1.7.
+For any five points in the plane, there exists a conic that passes through all of them.
+For a -scheme we define the associated exact scheme
 Accordingly, for a set we define
 where is a set of simple points.
-Proposition 1.13 ([5], lemma 2.1).
+Proposition 1.8 ([6], Lem.2.1).
 The problem is -correct, for some if and only if the problem is -solvable.
@@
 The direction “only if” is evident.
-For the direction “if”, in view of Proposition 1.10, it is enough to prove that there exists a (simple) node such that the problem where and is -solvable.
-For this, in view of -solvability of we need just an -fundamental polynomial of the simple node
-It is easily seen that the polynomial is the desired -fundamental polynomial of ∎
+For the direction “if”, in view of Proposition 1.5(i), it is enough to prove that there exists a (simple) point such that the problem where and is -solvable.
+For this, in view of the -solvability of we need just to have an -fundamental polynomial of the simple point
+Then, it is easily seen that the polynomial is the desired -fundamental polynomial of ∎
+Denote
+Corollary 1.9.
+Suppose that the problem is -solvable. Then we have that
+Proof.
+Remark 1.10 ([3], rem.1.14).
 Let and
-Remark 1.14 ([3], rem.1.14).
 If the problem is -solvable, then so is the problem with sufficiently close to
+If the problem is -solvable, with all belonging to an -dimensional linear subspace then so is the problem where the projection of on is
+Until the end of this subsection, we study the bivariate case.
+In this case we use the following simplified notation:
+In this case instead of (1.1) it is more convenient to use the notation
+Then the equality (1.5) and the second equality of (1.2) become:
+A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation .
+In particular, by we denote a linear polynomial from and the line defined by the equation Similarly, by we denote a quadratic polynomial from and the conic defined by the equation
 Definition 1.11.
-Let be a problem, where is given by (1.1), and be a curve of degree Then we define the scheme as follows
+Let be a problem, where is a set given by (1.3), and be a plane curve of degree Then we define the scheme as follows
 where if and otherwise.
-Denote by the number of nodes from on the curve counted with the multiplicities from
-Note that in what follows, this notation, as well as the previous definition, mainly will be used in cases where the curve is a line or a conic.
-We will need the following well-known proposition (see e.g. [2], lemma 4):
+Denote by the number of points from on the curve counted with the multiplicities from
+Note that this notation, as well as the previous definition, mainly will be used in the cases where the curve either is a line or a conic.
+For denote
+We will need the following well-known proposition:
 Proposition 1.12.
-Let be a scheme with and be a set of nodes with where is a line. Then
+Let be a scheme with and be a set of points with where is a line. Then
+where and
+Corollary 1.13.
+Suppose is a problem, where is -scheme. Suppose also is a line with Then the problem is -solvable if and only if the problem is -solvable, where
+1.2 Some results in the multivariate case
 Set
-The following proposition we will use for lines and (irreducible) conics.
+The following proposition will be used for lines and conics.
 Proposition 1.14.
-Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then for any curve of degree we have that
+Let be a scheme. If the problem is -solvable for a set of points then for any plane curve of degree we have that
 Proof.
-Let the problem be -solvable for a set of nodes Assume, by way of contradiction, that for a curve of degree we have that
+First consider the case Let the problem be -solvable for a set of points Assume, by way of contradiction, that for a curve of degree we have that
 Then, Proposition 1.8 implies that
@@
 Moreover, by making use of Remark 1.10 we may assume that the set contains no point from the curve
-Consider the following scheme
+Now, consider the following scheme
 The above condition on the curve in view of (1.10), means that
-Then let us check that is -scheme. Indeed, by using (1.7) and (1.9), we get
+Then let us check that is -scheme. Indeed, by using (1.10) and (1.12), we obtain
 Therefore, in view of Proposition 1.6, there exists a polynomial such that
-Now one readily obtains for the polynomial that
-This, in view of Proposition 1.10, contradicts the -correctness of the problem ∎
+Now we readily get that the polynomial satisfies the condition:
+Now consider the case of general dimension: Assume that does not satisfy the condition, i.e., there is a plane curve passing through more than of its points.
+Consider the respective subproblem restricted to the interpolation points belonging to the curve As we proved above, this subproblem is not -solvable in the case of Therefore the original problem also is not -solvable.
+∎
 By applying Proposition 1.14 for the cases we obtain the following two necessary conditions for -solvability:
 Corollary 1.15.
-Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then:
+Let be a scheme. If the problem is -solvable for a set of points then:
 For any line we have that
@@
 Similarly, the condition (ii) will be referred to as the -condition.
-Note, that by taking into account the above condition (i) for the lines passing through any two nodes, we obtain
+Definition 1.16.
+A scheme is called regular if there exists a set of points such that the problem is -independent.
+By taking into account the above condition (i) for the lines passing through any two points, and the condition (ii) for the conics passing through any five points, one readily obtains
 Corollary 1.17.
-Let be a scheme with non-increasing sequence of multiplicities. If the problem is solvable for a set of nodes then
-Now we are in a position to present the result (slightly reformulated), which we will extend.
-Theorem 1.20 ([2], th. 9).
+Let be a scheme with non-increasing sequence of multiplicities. If is regular then we have that
+Now we are in a position to present the result, which we are going to extend. Actually, it is slightly reformulated (see [2], 2nd par. after the proof of Thm.9).
+Theorem 1.18 ([2], Thm.9).
 Let be a scheme, satisfying the condition
-Then the problem is -solvable if and only if there is no line passing through more than nodes of
-This theorem can be regarded as a natural generalization of the following classical theorem of Severi (1921).
-Theorem 1.21 (Severi [7], th. 9).
-Let be a scheme. The problem is solvable for any node set if and only if
+Then the problem is -solvable if and only if there is no line passing through more than points of
+This theorem can be regarded as a natural extension of the following classical Theorem of Severi (1921).
+Theorem 1.19 (Severi [8], Thm.9).
+Let be a scheme. Then the problem is -solvable for any point set if and only if
+Until the end of the subsection, we study the bivariate case again.
 From Theorem 1.18 we get
 Corollary 1.20.
-For any scheme with non-increasing sequence of multiplicities, we have that
-Indeed, consider the problem where in no three nodes are collinear. Then according to Theorem 1.20 the problem is -solvable. Hence is -scheme.
+Suppose that a scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if only if
+Proof.
+For denote
+Next proposition is an extension of Proposition 1.12:
 Proposition 1.21.
 Let be a scheme with non-increasing sequence of multiplicities such that
-Suppose is a set of nodes with where is an irreducible conic. Then we have that
+Suppose is a set of points with where is an irreducible conic. Then we have that
+where and
 Proof.
@@
 Note that
-The above inequality, in view of irreducibility of implies that the -condition is satisfied.
-Thus, in view of Theorem (1.20), we obtain that the problem is -solvable.
-Similarly we obtain that the problem is -solvable.
-Now consider the following linear spaces:
-Note that In view of … we have that
+The above inequality, in view of irreducibility of implies that the -condition is satisfied for the problem .
+Thus, in view of Theorem (1.18), we obtain that is -solvable. Similarly we obtain that the problem is -solvable.
+Now, consider the following linear spaces:
+Note that Since the problems and are solvable, we get from Corollary 1.9, that
 Thus we obtain that ∎
@@
 Let be a scheme with non-increasing sequence of multiplicities such that
-Then we readily get that implies that While implies that
-Both these schemes are exact.
-According to the above proof the problems are also solvable, provided with any irreducible conic.
-The latter condition in this case means that no three nodes are collinear.
-Thus the problems are correct:
-This holds also with the scheme replaced by Therefore Proposition 1.23 is valid also in the case since the statement reduces to
+Then implies that While implies that
+Now, let Note that is exact.
+The proof of Proposition 1.21 implies that the problem is solvable, provided that with any irreducible conic.
+The latter condition in this case means that no three points of are collinear.
+Thus the problem is correct:
+Note that also (see the above proof) is correct, since if then Therefore Proposition 1.21 is valid also in the case and the statement (1.13), i.e., reduces to
 Remark 1.23.
-Note that according to Corollary 1.22 for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
+Note that, according to Corollary 1.20, for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
 In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme.
-Indeed, for a counterexample, we just add a fifth multiplicity to the schemes and of the above remark.
-This is why in the next theorem, unlike Theorem 1.20, we have the condition: is -scheme.
-2 The result
+Indeed, for a counterexample, we just add a fifth multiplicity to the scheme of the above remark.
+This is why, in the following theorem we have an additional condition, compared to Theorem 1.18: is -scheme.
+2 The main result
 In this section we will prove the following
@@
 Let be a -scheme satisfying the condition
-Then the problem is -solvable if and only if
-there is no line passing through more than nodes of ,
-there is no conic passing through more than nodes of
+Then the problem is -solvable if and only if:
+There is no line passing through more than points of
+There is no conic passing through more than points of
 Proof.
-The direction “only if” follows from Corollary 1.18. Let us prove the direction “if”.
-Note that the above condition (ii) means that there is no conic passing through all the nodes of
+The direction “only if” follows from Corollary 1.15. Let us prove the direction “if”. First we consider the (main) case
+Note that the above condition (ii) means that there is no conic passing through all the points of
 In view of Theorem 1.18 we can assume, in what follows, that
 From now on suppose also that is a scheme with non-increasing sequence of multiplicities:
-In view of Corollary 1.19 we have that
+By using the condition (i) for the line passing through we get
 We can assume also that
-Indeed, the only - scheme, in the case is the Taylor scheme which does not satisfy (2.2). While if then the conditions (2.2) and (2.4) imply that which is not - scheme.
+Indeed, assume that Then the only -scheme is the the Taylor scheme But the latter does not satisfy (2.1). While if then the conditions (2.1) and (2.3) imply that which is not -scheme.
 We will apply induction on the degree of the scheme
@@
 In the case there is only one -scheme satisfying the conditions (2.1) and (2.4).
-According to the condition (ii) of Theorem, the six simple nodes of the set do not belong to a conic.
-Since is exact scheme, by Proposition 1.5 (iii), we get that is -correct.
+According to the condition (ii) of Theorem, the six simple points of the set do not belong to a conic.
+Since is exact scheme, we obtain from Proposition 1.5(ii), that is -correct.
 Now, assuming the statement holds for all degrees we will prove it for where
-Step 2. Assume that there is a line which passes through nodes of i.e., the following condition is satisfied:
+Step 2. Assume that there is a line which passes through points of i.e., the following condition is satisfied:
 Consider the scheme (see (1.9))
 From (1.10), in view of the condition (2.5), we have that
-Now let us check that is - scheme. Indeed, by using (2.7), we get
-The condition (2.1) for readily follows from the fact that passes through at least two nodes of
-Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing nodes of Assume conversely that there is such a line Then evidently and, in view of the relation
-we obtain that contains nodes from Since contains nodes both from and we conclude that the line does not intersect at a node of Therefore, the (reducible) conic passes through nodes of which contradicts the condition (ii) of Theorem.
-Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing nodes of In view of (2.9) this means that passes through all the nodes with
-Note that if passes through at least three nodes of then in (2.9) we would obtain and hence the -condition is satisfied.
-It remains to consider the case where passes through exactly two nodes, say, In this case, in view of (2.6), we have that According to (2.5) we get that Hence we have that Thus passes through all the nodes of which is contradiction.
+Next, we want to apply the induction hypothesis for the problem For this let us check first that is -scheme. Indeed, by using (2.6), we get
+The condition (2.1) for readily follows from the fact that passes through at least two points of
+Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing points of Assume conversely that there is such a line Then evidently and, in view of the relation
+we obtain that contains points from Since contains points both from and we conclude that the line does not intersect at a point of Therefore, the (reducible) conic passes through points of which contradicts the condition (ii) of Theorem.
+Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing points of In view of (2.7) this means that passes through all the points with
+Note that if passes through at least three points of then in (2.7) we would obtain and hence the -condition is satisfied.
+It remains to consider the case where passes through exactly two points of , say, and In this case, in view of (2.5), we have that According to (2.4) we get that Hence we have that Thus passes through all the points of which is contradiction.
 Therefore, the problem is -solvable by virtue of the induction hypothesis.
-Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
+Now, Proposition 1.8 implies that there is a set consisting of simple points (see (1.7) and (1.8)), such that the problem
 where
-and the nodes of correspond to the last ones in
-Here, by making use of Remark 1.14, we can also assume that the set contains no node belonging to the line
+Here, by making use of Remark 1.10, we can also assume that the set contains no point belonging to the line
 Now consider the problem where
-Let us check first that is exact scheme. Indeed, in view of (2.7), we have that
-Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.10, (iii), that where and prove that
+Let us check first that is exact scheme. Indeed, in view of (2.6), we have
+Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.5(ii), that where and prove that
 By the choice of the set and the condition (2.5) we have that
-Now, Proposition 1.16 implies that where
-Now, in view of (2.10) and Proposition 1.10, we get and hence
+Then, Proposition 1.12 implies that where
+Now, in view of (2.8) and Proposition 1.5(ii), we get and hence
 Thus, the problem is -correct, which implies evidently that is -solvable. ∎
 From now on, in view of Step 2, we can assume that the -condition is satisfied for the scheme which implies that
-Step 3. Assume that there is a line passing through at least nodes of Consider the following scheme (see (1.6))
+Step 3. Assume that there is a line passing through at least points of Consider the following scheme (see (1.9))
 We have that
 On the other hand, in view of Step 2, the -condition holds. Thus according to Theorem 1.18 we obtain that the problem is -solvable.
-Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
+Now, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
 where
-Again, by making use of Rremark 1.14, we can assume also that the set contains no node belonging to the line
-Then denote by a set of simple nodes in
+Again, by making use of Remark 1.10, we can assume also that the set contains no point belonging to the line
+Then denote by a set of simple points in
 Consider the problem where and
 Let us check first that is exact scheme. Indeed, we have that
-Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
+Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
 In view of the choice of the sets and we have that
 Therefore, Proposition 1.12 implies that where
-Now, in view of (2.13) and Proposition 1.10, we get and hence
+Now, in view of (2.10) and Proposition 1.5(ii), we get and hence
 Thus, the problem is -correct, which implies that is -solvable. ∎
-From now on, in view of Step 3, we may assume that there are no three collinear nodes in the set
-Step 4. First note that the condition (ii) of Theorem, in view of Lemma 1.12, implies that
-Consider a conic passing through the nodes Since no three nodes are collinear the conic is irreducible.
+From now on, in view of Step 3, we may assume that there are no three collinear points in the set
+Step 4.
+First note that the condition (ii) of Theorem, in view of Corollary 1.7, implies that Then consider a conic passing through the points The conic is irreducible, since no three points are collinear.
 Note that may not satisfy the condition (2.2). Thus, to apply Theorem 1.18, let us verify that
 Indeed, if then and, in view of (2.9), this is evident.
-Now, assume conversely, that the inequality holds for some
-Then, in view of (2.3), we have that and therefore
+Now, assume conversely, that the inequality holds for some Then, in view of (2.2), we have that and therefore
 On the other hand we have that if Hence we have that which contradicts the condition (2.1).
-Thus, according to Theorem 1.20, the problem is -solvable.
-Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
+Thus, according to Theorem 1.18, the problem is -solvable. Therefore, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
 where
-By making use of Remark 1.14 we can assume that the set contains no node belonging to the conic
-Next, denote by a set of simple nodes in
+By making use of Remark 1.10 we can assume that the set contains no point belonging to the conic
+Next, denote by a set of simple points in
 Consider the problem where and
 Let us check first that is exact scheme. Indeed, we have that
-Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
+Now, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
 By taking into account the choice of the sets and we have that Therefore, Proposition 1.21 implies that where
-Now, in view of (2.15) and Proposition 1.10, we get and hence
+Then, in view of (2.11) and Proposition 1.5(ii), we get and hence
 Thus, the problem is -correct, which implies that is -solvable. ∎
+Step 5. It remains to consider the case of general dimension.
+Assume that is a set of points in dimension satisfying the and -conditions with respect to the problem Then we can project it to a set of points in dimension, with the same property.
+Continuing in this way we will obtain a set of points in a plane, with the same property.
+Now, the problem is -solvable, since Theorem has already been proven in the bivariate case.
+From here, by using times Remark 1.10, we obtain that the original problem is -solvable too.
+∎
+A counterexample.
+Suppose the scheme is the following Thus (2.12) takes place. Also we have that is -scheme:
+Suppose the corresponding set of points is chosen as follows: The points all belong to an irredusible conic The points and belong to a line which does not contain any other point from
+It is easily seen that the conditions (i) and (ii) of Theorem 2.1 are satisfied:
+Let us verify that the Hermite problem is not -solvable.
+Indeed, by applying Corollary 1.13 with respect to the line we get that the problem is -solvable if and only if the problem is -solvable, where and
+Now, on the conic we have points of Hence, in view of Corollary 1.15(ii), the problem is not -solvable.
+From Theorem 2.1 we get
+Corollary 2.2.
+Suppose that a -scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if and only if and
+Proof.
+Consider the following exact scheme This scheme satisfies the conditions of above corollary, with one exception: According to Corollary 1.17 the scheme is not regular.